# Sustainability Risks under Lotka-Volterra Dynamics

### v1

@@ 1 Introduction
The record-breaking heat in recent years, along with other extreme weather conditions worldwide, has not only warned us about the devastating effects of global warming but also revived our interest in studying sustainability risks on a broader scale. The Charter for the UCLA Sustainability Committee defines sustainability as: “the integration of environmental health, social equity and economic vitality to create thriving, healthy, diverse and resilient communities for this generation and generations to come [^1] See: sustain.ucla.edu/what-is-sustainability/.. The World Commission on Environment and Development defines sustainability as “meeting the needs of the present without compromising the ability of future generations to meet their own needs.” (WCED, 1987). Subsequently, the UN advocates the Sustainable Development Goals (SDGs), which entail 17 specific goals that look at issues such as poverty, inequality, climate change, environmental degradation, peace and justice. The Conference of the Parties (COP) was initiated by the UNFCCC [^2] United Nations Framework Convention on Climate Change, established in 1992 at the Rio Earth Summit in Brazil, to address climate issues with a particular focus on emissions reduction. Since UNFCCC initiated the first COP [^3] Conference of the Parties meeting held in Berlin in 1995, there have already been 30 COP meetings held up to 2025. Over the past three decades, the COP meeting has evolved from its early function of establishing frameworks and protocols to its recent advocacy of specific actions, such as climate finance, carbon trading, and net-zero targeting.
@@
In finance, one of the most significant developments to address sustainability risks since the 1992 Rio Earth Summit is the introduction of the environmental, social, and governance (ESG) metric, which strives to establish a standard for disclosing sustainability-related information of public companies. Although the discussion of corporate social responsibility has a long history, ESG was not formally introduced and adopted until 2004, when the term ESG was coined in the UN-backed report Who Cares Wins (United Nations Global Compact, 2004). There is an increasing adoption of ESG reports by investors and an exponential growth of companies that start to provide ESG reporting, as well as agencies that offer ESG rating services. The dissemination of ESG has also led to an explosion of academic research on ESG in the field of corporate finance (Ioannou and Serafeim, 2019; Berg et al., 2022; Cheng et al., 2014; Dimson et al., 2015; Amel-Zadeh and Serafeim, 2018; Viehs et al., 2014; Friede et al., 2015; Pedersen et al., 2021; Fulton et al., 2013). By August 18, 2025, a simple search with the term ESG on sciencedirect yields more than 15,000 results [^4] See: https://www.sciencedirect.com/search?qs=ESG.
However, there is little evidence that the climate issue has been alleviated over the last two decades since the introduction and adoption of ESG. According to Fig. 2, there has been a steady rise in global temperature since 2000, which renders the years 2023 and 2024 the hottest two consecutive years in history (Xie et al., 2025). A recent CNBC news article reported that there are worries from the insurance industry that the world could soon become uninsurable [^5] Sam Meredith, Why insurers worry the world could soon become uninsurable, CNBC, 8 August, 2025.
https://www.cnbc.com/2025/08/08/climate-insurers-are-worried-the-world-could-soon-become-uninsurable-.html
due to the rising global temperature. Both practitioners and academics raise questions regarding the efficacy of ESG in addressing climate issues. Some point to a lack of standardisation and rating divergence among the rating agencies, while others express their concerns about the scandals involving some high-profile public companies over recent years (Johnson, 2025). In academic research, most studies focus on trivial issues such as company performance, equity valuation, and investor trust, yet overlook the long-term efficacy of ESG in measuring long-term sustainability risks.
@@
The dire situation of climate change, however, reveals only a tip of the iceberg of sustainability risks resulting from human activities. It is not only the status quo, but also the avant-garde technologies that would pose significant sustainability risks, potentially leading to destructive environmental and social consequences. As a pinnacle of disruptive technology of this era, the widespread adoption of AI may lead to unintended environmental and social risks arising from a substantial rise in energy consumption and the displacement of traditional employment and social ethics (Torres, 2025). The production of electronic gadgets in today’s digital era may lead to severe water and soil pollution, as the production of such products requires a high consumption of lithium and rare earth materials that can have an adverse impact on the environment (Duporte, 2022). The adoption of electric vehicles may not reduce carbon emissions if the electricity is generated from coal instead of clean sources (Zeng et al., 2021). Sustainability risk is systemic, whereby the failure of one part may lead to the collapse of the entire system (Schweizer and Juhola, 2024). While ESG can undoubtedly provide a guiding framework for individual companies to act responsibly and sustainably at an individualist level, we also need to understand, from a systemic perspective, whether the actions taken can have a real, positive impact on the environment and society, and how sustainability can be achieved through collective efforts.
@@
Todorov and Marinova (2011) conduct a comprehensive survey of sustainability models. They base their analysis on the three pillars of sustainability (environment, economy, humanity) and categorise the sustainable development model into five groups: pictorial visualisation models, quantitative models, physical models, conceptual models, and standardising models. Within the last category of standardising model, the most well-known would be the landmark study “The Limits to Growth” (Meadows et al., 1972). The book is a commissioned work by the Club of Rome, which had led to the creation of the World3 model by a team of MIT scientists: a large-scale computer model that simulates the long-term ramifications of unconstrained exponential economic growth by examining population, industrial output, food production, pollution, and consumption of non-renewable resources (citation). The World3 model has been recalibrated and updated by Meadows et al. (1992), Meadows et al. (2004), and more recently by Herrington (2021). The original World3 model and its later variants generally assert that the global system will reach its limits within a foreseeable period in a business-as-usual (BAU) scenario. Yet, humans have the capacity to avert the doomsday situation if preventive measures are taken.
@@
Since the seminal work of Meadows et al. (1972), the World3 model has been the subject of considerable controversy and criticism, particularly following the first book’s failure to predict resource depletion by 1990. Numerous studies argue that the World3 model overlooks the forces of the market and innovation: the increasing scarcity of resources would raise their prices, forcing companies to find alternative resources in the short term and to innovate for other technological solutions in the long term (Kaysen, 1972; Solow, 1973). From a modelling perspective, several studies point out that the World3 model is sensitive to parameter changes and initial conditions (Castro, 2012). Indeed, the World3 model is a high-order, non-linear dynamic system with 12 state variables and internal delay functions with complex and intertwined feedback loops [^6] See: https://github.com/cvanwynsberghe/pyworld3., which renders it theoretically challenging to derive its closed-form analytical solution. It is widely recognised that the predictive horizon of complex models is constrained by their degree of complexity, particularly for the case of a chaotic system (Lorenz, 1963). It is questionable whether the very long-term trend projection derived from such a complex model is valid over the very long time horizon. On the contrary, a dynamic model with a more parsimonious structure may offer better predictive power over the short to medium term, which is vital for setting the stage for policy debates and guiding regulatory measures over a shorter time horizon. Furthermore, the market economy is a dynamic process constantly fueled by disruptive innovation and incessant entrepreneurial pursuits. Hence, the model should capture the sustainability risk resulting from the creative destruction process in Schumpeterian terms (Schumpeter, 1934). Despite its controversy, the World3 model has paved the way for researchers later on to study sustainability risks from a systems dynamics perspective.
@@ 2 Model Set-up
While the three state variables depict a generalised situation, they can also be decomposed by the sub-factors, where Euler’s theorem can aggregate the three state variables, following the common practice in financial risk management literature:
$E$ $=$ $\sum_{i=1}^{n_{1}}\alpha^{E}_{i}E_{i}=\sum_{i=1}^{n_{1}}\frac{\partial E}{\partial E_{i}}E_{i},$ (5)
$V$ $=$ $\sum_{i=1}^{n_{2}}\alpha^{V}_{i}V_{i}=\sum_{i=1}^{n_{2}}\frac{\partial V}{\partial V_{i}}V_{i},$ (6)
$S$ $=$ $\sum_{i=1}^{n_{3}}\alpha^{S}_{i}S_{i}=\sum_{i=1}^{n_{3}}\frac{\partial S}{\partial S_{i}}S_{i},$ (7)
$N$ $=$ $n_{1}+n_{2}+n_{3},$ (8)
where $E$, $V$ and $S$ consist $n_{1}$, $n_{2}$ and $n_{3}$ number of sub-factors summing up to $N$. $\alpha^{E}_{i}=\frac{\partial E}{\partial E_{i}}$, $\alpha^{V}_{i}=\frac{\partial V}{\partial V_{i}}$, $\alpha^{S}_{i}=\frac{\partial S}{\partial S_{i}}$ are the three Euler weights.
@@ 2.1 The definition of the three state variables
The World Commission on Environment and Development (WCED), established by the UN and led by former Norwegian Prime Minister Gro Brundtland, defines sustainable development as meeting the needs of the present without compromising the ability of future generations to meet their own needs (WCED, 1987). We may translate this definition as the three state variables in a dynamical system that co-exist with, as well as are interdependent upon each other, so that none will become suppressed to extinction due to the overexpansion of the other(s). Mathematically, this can be expressed as:
 $\exists\text{ }T_{1}\text{ such that when }T>T_{1},\ E>0,V>0,S>0\text{.}$
The measurement of the three state variables, however, is by no means a simple task. The commonly adopted measures of economic, social, and environmental indicators often have their limitations, and sometimes they overlap with one another in certain sub-dimensions. First, it is challenging to find measures of economic development, since existing economic measures focus on growth rather than development. Second, many social development indicators are intertwined with economic and environmental measures, thus deserve further specification in our context. Third, environmental measurement should take into account not only the knowns, but also the unknowns, such as the new environmental challenges arising from the new era of AI, energy transition, and digitalisation.
@@ 2.1.1 Measuring economic development
The method by which economic development is measured deserves further discussion. We begin with the assertion that economic development measures, from a Schumpeterian perspective of economic development, should reflect the progress of technology and innovation, whether it is the existing technology (the status quo) or the emerging new technology (the avant-garde). One may be tempted to adopt simple macroeconomic measures such as GDP, which is defined as “the total market value of all final goods and services produced within a country in a given period of time” (Black et al., 2009). We argue that it is inappropriate to take this over-simplistic approach for two reasons. First, GDP represents a flow, rather than a stock, of a given period. The LV model uses the stock of each state variables instead of their flow. The delayed impact of GDP on current social and environmental factors should also be considered. A deeper issue concerns the distinction between economic growth and economic development. While growth can be measured in standardised and quantitative terms, economic development involves more profound structural changes beyond quantitative measures. J.A. Schumpeter’s Theory of Economic Development depicts the vital role of entrepreneurs in the process of economic development, as they constantly seek profit by innovation, resulting in the creative destruction [^8] The term Creative Destruction was not coined until Schumpeter’s later works Capitalism, Socialism, and Democracy (Schumpeter, 1942). and structural changes that serves as the ultimate engine of economic development (Schumpeter, 1934). Hence, innovation and entrepreneurship should play a vital role in measuring economic development from a Schumpeterian perspective. The Global Innovation Index (GII), launched in 2007 at WIPO [^9] World Intellectual Property Organisation, provides a comprehensive framework in measuring innovation (Dutta and Caulkin, 2007; WIPO, 2024), which can potentially be adopted in our context.
@@ 2.1.2 Measuring social development
Yet the CSS measure of social development overlooks inequality as a key issue of social development. The earliest and perhaps most well-known inequality measure is the Gini index, proposed by Gini (1912), where inequality is measured on a scale from 0 (perfect equality) to $1$ (perfect inequality). More recently, Aspalter (2023) systematically examines the causes of super inequality from his theory of Z-efficiency that addresses a broader scope of social, behavioural, and institutional inefficiency. The $Z$-efficiency theory is extended from Leibenstein’s $X$-efficiency theory of Leibenstein (1966), and it encompasses a set of multi-dimensional social efficiency measures, including $M$-efficiency (removal of managerial barriers), $N$-efficiency (removal of environmental and physical barriers), and $Y$ & $Z$-efficiency (removal of personal-level barriers). The $Z$-efficiency theory can serve as an important measure of social development in our context.
@@ 3.4 The merged full model and the feedback loops
4 The full 3D EVS system and the sustainable scenario
@@ 4.1 The baseline model
We merge the EV, ES, VS subdynamics to form the baseline 3D EVS model in Eq. 22 - 24.
@@
4.2 A formal definition of sustainability
Figure 5: A formal definition of sustainability, reproduced based on Vandermeer (1990)
We define sustainability in the EVS system as the condition where all dimensions remain in the positive octant and exhibit resilience away from extinction. Following Vandermeer (1990), the sustainability constraints can be formally derived by investigating the time derivative of each state variable at the near extinction level with reference to its complementary subdynamics. For example, in the case of an ES subsystem, we may set $V=0$ to derive the fixed point at the ES subsystem $(E^{\star},S^{\star})$ and set $\frac{dV}{dt}>0$ to derive its sustainability condition. We may interpret this equation in a hypothetical situation of near environmental depletion. With the injection of a tiny environmental aid, the sign of $\frac{dV}{dt}$ would determine whether the ecological system will recover on its own or not. The other two state variables can be examined in the same manner, as illustrated in Fig. 5. Formally, we define that:
@@ 4.2 A formal definition of sustainability
$\frac{dE}{dt}\bigg{|}_{(V,S)\to(V^{\star},S^{\star})}$ $>$ $0,$ (27)
$\frac{dV}{dt}\bigg{|}_{(E,S)\to(E^{\star},S^{\star})}$ $>$ $0,$ (28)
$\frac{dS}{dt}\bigg{|}_{(E,V)\to(E^{\star},V^{\star})}$ $>$ $0,$ (29)
where $(V^{*},S^{*})$, $(E^{*},S^{*})$, and $(E^{*},V^{*})$ are the three fixed points at the extinction level of the complementary dimension. Similarly, the bearable, equitable, and viable paradigm is the choice of V&S, E&S, and E&V while sacrificing the remaining dimension. Mathematically, we may define:
Definition 1.1 (Bearability): An EVS system is bearable when
$\frac{dE}{dt}\bigg{|}_{(V,S)\to(V^{*},S^{*})}<0,\text{{ ceteris paribus, w.r.t. Definition 1.}}$ (30)
Definition 1.2 (Equitability): An EVS system is equitable when
$\frac{dV}{dt}\bigg{|}_{(E,S)\to(E^{*},S^{*})}<0,\text{{ ceteris paribus, w.r.t. Definition 1.}}$ (31)
Definition 1.3 (Viability): An EVS system is viable when
$\frac{dS}{dt}\bigg{|}_{(E,V)\to(E^{\star},V^{\star})}<0,\text{{ ceteris paribus, w.r.t. Definition 1.}}$ (32)
The EVS system set by Eq. 22 - 24 takes an explicit form of a 3D Lotka-Volterra model. The parameter space of the sustainable scenario can simply be derived with regard to Definition 1.
Proposition 1 (Sustainability Boundary): The EVS system of Eq. 22 - 24 is sustainable when the following conditions set by Eq. 33 - 35 are fulfilled:
@@
[…]
Proof: provided in A.
We augment this proposition with a critical economic interpretation: the trajectory of motion under a particular parameter set represents a bottom-up outcome. In contrast, the parameter set itself represents a top-down precondition that can be altered through public intervention. Sustainability risks may arise not only from existing technology (as epitomised by the consumption of fuels), but also from the unpredictable, disruptive innovation (the Schumpeterian part). The emergence of sustainability risks is characterised by the violation of the aforementioned sustainability condition within a particular parameter set. The public sector, however, can alter the parameter set through various forms of intervention, such as fiscal stimulus, environmental regulation, and mission-oriented innovation policies to ensure that the parameter set remains within the sustainable regime (the Keynesian part). In other words, the conditions set by Eq. 33 - 35 would determine whether the laissez-faire market force would fail to deliver the desired sustainable development outcome, and whether public intervention is needed to ensure sustainability. This interpretation aligns well with the Keynesian-Schumpeter synthesis discussed by Dosi et al. (2010), as well as the concept of the entrepreneurial state and mission-oriented innovation policies proposed by Mazzucato (2013).

### v2

@@ 1 Introduction
The record-breaking heat in recent years, along with other extreme weather conditions worldwide, has not only warned us about the devastating effects of global warming but also revived our interest in studying sustainability risks on a broader scale. The Charter for the UCLA Sustainability Committee defines sustainability as: “the integration of environmental health, social equity and economic vitality to create thriving, healthy, diverse and resilient communities for this generation and generations to come (UCLA Sustainability, 2025). The World Commission on Environment and Development defines sustainability as “meeting the needs of the present without compromising the ability of future generations to meet their own needs.” (WCED, 1987). Subsequently, the UN advocates the Sustainable Development Goals (SDGs), which entail 17 specific goals that look at issues such as poverty, inequality, climate change, environmental degradation, peace and justice. The Conference of the Parties (COP) was initiated by the UNFCCC [^1] United Nations Framework Convention on Climate Change, established in 1992 at the Rio Earth Summit in Brazil, to address climate issues with a particular focus on emissions reduction. Since UNFCCC initiated the first COP [^2] Conference of the Parties meeting held in Berlin in 1995, there have already been 30 COP meetings held up to 2025. Over the past three decades, the COP meeting has evolved from its early function of establishing frameworks and protocols to its recent advocacy of specific actions, such as climate finance, carbon trading, and net-zero targeting.
@@
In finance, one of the most significant developments to address sustainability risks since the 1992 Rio Earth Summit is the introduction of the environmental, social, and governance (ESG) metric, which strives to establish a standard for disclosing sustainability-related information of public companies. Although the discussion of corporate social responsibility has a long history, ESG was not formally introduced and adopted until 2004, when the term ESG was coined in the UN-backed report Who Cares Wins (United Nations Global Compact, 2004). There is an increasing adoption of ESG reports by investors and an exponential growth of companies that start to provide ESG reporting, as well as agencies that offer ESG rating services. The dissemination of ESG has also led to an explosion of academic research on ESG in the field of corporate finance (Ioannou and Serafeim, 2019; Berg et al., 2022; Cheng et al., 2014; Dimson et al., 2015; Amel-Zadeh and Serafeim, 2018; Viehs et al., 2014; Friede et al., 2015; Pedersen et al., 2021; Fulton et al., 2013). By 25 October, 2025, a simple search with the term ESG on sciencedirect yields more than 17,000 results [^3] See: https://www.sciencedirect.com/search?qs=ESG for the latest updates.!
However, there is little evidence that the climate issue has been alleviated over the last two decades since the introduction and adoption of ESG. According to Fig. 2, there has been a steady rise in global temperature since 2000, which renders the years 2023 and 2024 the hottest two consecutive years in history (Xie et al., 2025). A recent CNBC news article reported that there are worries from the insurance industry that the world could soon become uninsurable due to the rising global temperature (CNBC, 2025). Both practitioners and academics raise questions regarding the efficacy of ESG in addressing climate issues. Some point to a lack of standardisation and rating divergence among the rating agencies, while others express their concerns about the scandals involving some high-profile public companies over recent years (Johnson, 2025). In academic research, most studies focus on trivial issues such as company performance, equity valuation, and investor trust, yet overlook the long-term efficacy of ESG in measuring long-term sustainability risks.
@@
The dire situation of climate change today can be seen as a revelation of a more profound issue: while technological progress serves as a fundamental force of economic growth and human development, its unintended consequences can also lead to devastating sustainability risks that demand early intervention. Furthermore, it is not only the status quo, but also the avant-garde technologies that would pose significant sustainability risks with profound social and ecological consequences. While global warming is a culmination of decades-long consumption of fossil fuels since the advent of the internal combustion engine (ICE) over a hundred years ago, the disruptive technology of the digital era would potentially lead to an even greater spillover effect with a shortened time horizon of impact, both in positive and negative terms. The widespread adoption of AI may lead to a substantial rise in energy consumption and the displacement of traditional employment and social ethics (Torres, 2025). The production of electronic gadgets may lead to severe water and soil pollution, as the production of such products requires a high consumption of lithium and rare earth materials that can have an adverse impact on the environment (Duporte, 2022). The adoption of electric vehicles may not reduce carbon emissions if the electricity is generated from coal instead of clean sources (Zeng et al., 2021). Sustainability risk is systemic, whereby the failure of one part may lead to the collapse of the entire system (Schweizer and Juhola, 2024). Therefore, it is vital to design a mission-oriented [^4] The term mission-oriented innovation policy is coined by Mazzucato (2013) in describing the critical role of the public sector in promoting innovation beyond the profit incentives of the private sector. policy framework that not only encourages innovation but also guides R&D spending to counterbalance the adverse effects of disruptive technology, with the ultimate goal of achieving a balanced growth trajectory that aligns with the UN’s Sustainable Development Goals.
@@
Todorov and Marinova (2011) conduct a comprehensive survey of sustainability models. They base their analysis on the three pillars of sustainability (environment, economy, humanity) and categorise the sustainable development model into five groups: pictorial visualisation models, quantitative models, physical models, conceptual models, and standardising models. Within the last category of standardising model, the most well-known would be the landmark study “The Limits to Growth” (Meadows et al., 1972). The book is a commissioned work by the Club of Rome, which had led to the creation of the World3 model by a team of MIT scientists: a large-scale computer model that simulates the long-term ramifications of unconstrained exponential economic growth by examining population, industrial output, food production, pollution, and consumption of non-renewable resources. The World3 model has been recalibrated and updated by Meadows et al. (1992), Meadows et al. (2004), and more recently by Herrington (2021). The original World3 model and its later variants generally assert that the global system will reach its limits within a foreseeable period in a business-as-usual (BAU) scenario. Yet, humans have the capacity to avert the doomsday situation if preventive measures are taken.
@@
Since the seminal work of Meadows et al. (1972), the World3 model has been the subject of considerable controversy and criticism. Numerous studies argue that the World3 model overlooks the forces of the market and innovation: the increasing scarcity of resources would raise their prices, forcing companies to find alternative resources in the short term and to innovate for other technological solutions in the long term (Kaysen, 1972; Solow, 1973). From a modelling perspective, several studies point out that the World3 model is sensitive to parameter changes and initial conditions (Castro, 2012). Indeed, the World3 model is a high-order, non-linear dynamic system with 12 state variables and internal delay functions with complex and intertwined feedback loops, which renders it theoretically challenging to derive its closed-form analytical solution (Vanwynsberghe, 2021). It is widely recognised that the predictive horizon of complex models is constrained by their degree of complexity, particularly for the case of a chaotic system (Lorenz, 1963). It is questionable whether the very long-term trend projection derived from such a complex model is valid over the very long time horizon. On the contrary, a dynamic model with a more parsimonious structure may offer better predictive power over the short to medium term, which is vital for setting the stage for policy debates and guiding regulatory measures over a shorter time horizon. Furthermore, the market economy is a dynamic process constantly fueled by disruptive innovation and incessant entrepreneurial pursuits. Hence, the model should capture the sustainability risk resulting from the creative destruction process in Schumpeterian terms (Schumpeter, 1934). Despite its controversy, the World3 model has paved the way for researchers later on to study sustainability risks from a systems dynamics perspective.
@@ 2.1 The definition of the three state variables
This section discusses the measurement and the components of the three state variables, which is by no means a simple task. The commonly adopted measures of economic, social, and environmental indicators often have their limitations, and sometimes they overlap with one another in certain sub-dimensions. It is necessary to provide a taxonomy to facilitate further analysis regarding their competitive and cooperative relationships in the Lotka-Volterra setting.
@@ 2.1.1 Measuring economic development
The letter E represents not only the economy, but also entrepreneurship. From a Schumpetarian perspective, such measures should be able to characterise economic development contributed by entrepreneurship and innovation. It also should reflect the economic impacts arising from the continuous accumulation of knowledge, either proprietary knowledge in the form of various kinds of IP such as patents or copyrights, or common, open source knowledge that has real impacts on productivity and economic growth. Both the existing knowledge (the status quo) and the emerging new knowledge (the avant-garde) would contribute to economic development, with varying degrees of impact over their lifetime. One may be tempted to adopt simple macroeconomic measures such as GDP, which is defined as “the total market value of all final goods and services produced within a country in a given period of time” (Black et al., 2009). We argue that it is inappropriate to take this oversimplistic approach for two reasons. First, GDP represents a flow, rather than a stock, of a given period, which renders it incomparable with other environmental and social indicators that are stocks in nature. A deeper issue concerns the distinction between economic growth and economic development. While growth can be measured in standardised and quantitative terms, economic development involves more profound structural changes beyond quantitative measures. J.A. Schumpeter’s Theory of Economic Development depicts the vital role of entrepreneurs in the process of economic development, as they constantly seek profit by innovation, resulting in the creative destruction [^6] The term Creative Destruction was not coined until Schumpeter’s later works Capitalism, Socialism, and Democracy (Schumpeter, 1942). and structural changes that serves as the ultimate engine of economic development (Schumpeter, 1934). It is therefore necessary to integrate innovation-oriented measures such as the Global Innovation Index [^7] The Global Innovation Index (GII), launched in 2007 by the World Intellectual Property Organisation (WIPO), provides a comprehensive framework in measuring innovation on a worldwide basis (Dutta and Caulkin, 2007; WIPO, 2024). (GII) as a measurement of economic development. The innovation-led productivity growth, however, is not a guarantee for sustainable development. The Jevons Paradox is the observation that improvements in resource efficiency can lead to a rising demand of that resource, thus posing sustainability risks (Jevons, 1865). Therefore, it is equally important to measure social development and environmental quality to gain a comprehensive understanding of sustainability risks arising from innovation, which will be further elaborated in the following sections.
@@ 2.1.2 Measuring social development
Yet the CSS measure of social development overlooks inequality as a key issue of social development. The earliest and perhaps most well-known inequality measure is the Gini index, proposed by Gini (1912), where inequality is measured on a scale from 0 (perfect equality) to $1$ (perfect inequality). More recently, Aspalter (2023) systematically examines the causes of super inequality from his theory of Z-efficiency that addresses a broader scope of social, behavioural, and institutional inefficiency. The $Z$-efficiency theory is extended from Leibenstein’s $X$-efficiency theory of Leibenstein (1966), and it encompasses a set of multi-dimensional social efficiency measures, including $M$-efficiency (removal of managerial barriers), $N$-efficiency (removal of environmental and physical barriers), $Y$-efficiency (removal of personal-level barriers), and $Z$-efficiency (removal of social and cultural barriers). The $Z$-efficiency theory can serve as an important measure of social development in our context.
@@ 3.4 The merged full model and the feedback loops
4 The full 3D EVS system
@@ 4.1 The baseline model
We merge the EV, ES, VS subdynamics to form the baseline 3D EVS model in Eq. 18 - 20.
@@
4.2 The sustainability and resilience conditions of the EVS system
Figure 5: An illustration of the sustainability and resilience conditions
We take a further step to investigate how sustainability can be modelled in the EVS system. Recall that WCED [^8] The World Commission on Environment and Development, established by the UN and led by former Norwegian Prime Minister Gro Brundtland defines sustainable development as meeting the needs of the present without compromising the ability of future generations to meet their own needs (WCED, 1987). We may translate this definition as the three state variables in the EVS system that coexist with, and are interdependent upon each other, so that none will become suppressed to extinction due to the overexpansion of the other(s). Mathematically, this can be formally defined as:
@@ 4.2 The sustainability and resilience conditions of the EVS system
 $\exists\text{ }T_{1}\text{ such that when }T>T_{1},\ E>0,V>0,S>0\text{.}$
In the current academic literature, the terms sustainability and resilience are often used interchangeably. However, the two concepts may have substantial differences in the practice of sustainable development and risk management. Roostaie et al. (2019) conduct a meta-analysis on the definitions of, and relationship between sustainability and resilience. They point out that sustainability can simply mean “to maintain the status quo and to not disappear”, while resilience refers to “the responses to shocks, surprises, unforeseen or hazardous disturbances”, thereby “having the capacity to persist in the face of change, to continue to develop with ever-changing environments”. In the context of the EVS system, resilience implies that the system can recover on its own in the event of near-extinction or near-destruction due to external shocks. Mathematically, we may define resilience as:
Definition 2 (Resilience): An EVS system is resilient when
@@
$\frac{dE}{dt}\bigg|_{E\to 0}$ $>$ $0,$ (23)
$\frac{dV}{dt}\bigg|_{V\to 0}$ $>$ $0,$ (24)
$\frac{dS}{dt}\bigg|_{S\to 0}$ $>$ $0.$ (25)
It is difficult to derive a general condition for global sustainability under the EVS system specified by Eq. 18 - 20, since the closed orbit from the EV subdynamics does not have an exact functional expression. However, it is possible to derive the sustainability condition at the fixed point of the subdynamics to determine whether it exhibits the resilience condition or not, following Vandermeer (1990). For example, in the case of an EV subsystem, we may set $S=0$ to find the fixed point at the EV subsystem $(E^{\star},V^{\star})$ and set $\frac{dS}{dt}>0$ to further derive its resilience condition. We may interpret this equation in a hypothetical situation of near-complete social destruction. With the injection of a tiny institutional aid, the sign of $\frac{dS}{dt}$ would determine whether the social system will recover on its own or not. The other two state variables can be examined in the same manner, as illustrated in Fig. 5. Formally, we define that:
Proposition 1 (Fixed Point Resilience): The fixed point $(E^{\star},V^{\star})$ of the EV sub-system defined by Eq. 18 - 20 is resilient on the social dimension when the following condition set by Eq. 26 is fulfilled:
@@
Proof: we first set $S=0$, rendering the system an EV subdynamics:
which yields
$\left\{\begin{aligned} E^{*}&=-\frac{r_{V}}{a_{21}},\\ V^{*}&=-\frac{r_{E}}{a_{12}}.\end{aligned}\right$. (28)
We can see that $E^{\star}>0$ and $V^{\star}>0$, since $r_{E}<0$, $a_{12}>0$, $r_{V}>0$, and $a_{21}<0$. Furthermore, the social resilience condition implies that:
$\frac{dS}{dt}=S(r_{S}+a_{31}E+a_{32}V)>0.$ (29)
Hence,
$r_{S}>\frac{a_{31}r_{V}}{a_{21}}+\frac{a_{32}r_{E}}{a_{12}}=\frac{a_{31}a_{12}r_{V}+a_{32}a_{21}r_{E}}{a_{21}a_{12}}.$ (30)
Since we assume that $a_{21}<0,a_{12}>0,a_{32}>0,r_{S}<0$, and $r_{E}<0$, it implies that:
$a_{21}(a_{12}r_{S}-a_{32}r_{E})$ $<$ $a_{31}a_{12}r_{V},$ (31)
which leads to:
@@
Similarly, it is also possible to derive the resilience conditions for the fixed points of the ES and VS system. However, the fixed points may be unattainable due to the constraints we set on the signs of the $r_{k}$ and $a_{ij}$ parameters, rendering some values of the fixed points negative. We put the mathematical derivation of the fixed points for the ES and VS system, along with its resilience conditions in A.
We augment Definition 1, Definition 2 and Proposition 1 with a critical economic interpretation: the trajectory of motion under a particular parameter set represents a bottom-up outcome. In contrast, the parameter set itself represents a top-down precondition that can be altered through public intervention. Sustainability risks may arise not only from existing technology (as epitomised by the consumption of fuels), but also from the unpredictable, disruptive innovation (the Schumpeterian part). The emergence of sustainability risks is characterised by the violation of the aforementioned sustainability condition within a particular parameter set. The public sector, however, can alter the parameter set through various forms of intervention, such as fiscal stimulus, environmental regulation, and mission-oriented innovation policies to ensure that the parameter set remains within the sustainable regime (the Keynesian part). In other words, certain parameter sets would determine whether the laissez-faire market force would fail to deliver the desired sustainable development outcome, and whether public intervention is needed to ensure sustainability. This interpretation aligns well with the Keynesian-Schumpeter synthesis discussed by Dosi et al. (2010), as well as the concept of the entrepreneurial state and mission-oriented innovation policies proposed by Mazzucato (2013).